OpenStudy (anonymous):

Solve –2x2 – 16x – 44 = 0

OpenStudy (uri):

@Sdowney Factorize or Quad?

OpenStudy (anonymous):

Quad

OpenStudy (uri):

http://blogs.discovermagazine.com/loom/files/2008/07/quadratic.jpg Use quad formula ^ a:-2 b:-16 c:-44

OpenStudy (uri):

Can you do it @Sdowney?

OpenStudy (anonymous):

\[ –2x^2 – 16x – 44 = 0 \rightarrow -2(x^2 + 8x + 22) = 0\] i.e.\[ x^2 + 8x + 22 = 0 \rightarrow \] Next use quadratif formula to find its solution

OpenStudy (anonymous):

Quadratic Formula?

OpenStudy (uri):

Plug in the values.

OpenStudy (anonymous):

Eww >S< Tattoo

OpenStudy (anonymous):

I'll try

OpenStudy (uri):

Al right.

OpenStudy (anonymous):

x=-b+-√b^2-4ac/2a x=16+-√24

OpenStudy (anonymous):

Now what?

OpenStudy (uri):

Hm wait.

OpenStudy (anonymous):

x=-4+-√24

OpenStudy (uri):

Did you get how @dpasingh explained you find the common from this eq: –2x2 – 16x – 44 = 0 -2 is the common so x^2+8x+22=0 Put the quad formula a:1 b:8 c:22 So,the answer is -8 + sqrt -72/2 -

OpenStudy (anonymous):

Hm. I found the answer. Thanks

OpenStudy (uri):

You're welcome :)